# Towards Class-wise Robustness Analysis

### v1

@@ 1 Introduction
Convolutional neural networks (CNNs) have achieved widespread success in various vision applications including image classification [11, 16, 31], image segmentation[19], and object detection [25]. Nonetheless, the existence of adversarial examples [7, 18, 6, 1] and common corruptions [13] like blurring, zooming, or Gaussian noise, poses challenges in their real-world deployment. Extensive efforts have been devoted to defending against these adversarial attacks and enhancing model generalization [30]. Adversarial training has emerged as a prominent defense technique to improve the robustness of classification models [2, 7]. Prior works have analyzed adversarial training from different perspectives including robust optimization [27], robust generalization [24], training strategy [32, 23, 28] and neural architecture [9, 8, 20, 10]. Howbeit, all these previous works have concentrated on improving the overall model robustness, neglecting the discrepancies in the robustness of individual classes. This imbalance in class-wise robustness can be exploited by attackers, who may target less robust classes. Therefore, a comprehensive understanding of adversarial training on class-wise robustness is crucial for improving the robustness of classification models in a meaningful way.
Recently a few studies have emphasized class-wise robustness disparity in adversarial training [26, 3]. However, their focus has been limited to comparing class-wise robust accuracy deviations to identify the vulnerable classes. While this is important, it is equally crucial to analyze class-to-class biases to gain insights into the latent space of robust models. Specifically, understanding which class labels are assigned erroneously or which classes are predominantly confused is essential. Therefore, we conduct a comprehensive study of class-wise robust accuracies with particular emphasis on false positives in class-wise misclassifications, to improve the understanding of class-wise biases.
@@ 2 Background: Network Evaluation Methods
When introducing new models for image classification tasks, one usually considers the network’s performance in terms of accuracy[11, 31, 16]. Such evaluation is most important, as we do not need a network that classifies randomly and wrongly. Additionally, the evaluation of the network’s robustness received more popularity in the last few years. Hence, a variety of robustness measures has been proposed [15]. On the one hand, common corruptions [13] have been introduced, which incorporate natural and system noise that can lead to misclassifications in the classification systems. On the other hand, adversarial attacks gained a lot of popularity to evaluate the network’s vulnerabilities. In consequence, a variety of attacks, e.g. [7, 18, 6] along with their defenses have been proposed [7]. The robust accuracy of a model is thereby usually defined as the model’s accuracy under a specific adversarial attack or corruption. Thus, most of these studies focus on improving the overall robust accuracy of models under attacks or when facing corruptions. A few recent works further investigated the class biases in model accuracy and robustness, arguing for a fair training process that allows classifying all classes about equally well [29, 26]. These works also showed that adversarial training seems to amplify class-wise biases in model accuracy. Yet, only little effort has been devoted to studying which classes pre-dominantly attract incorrectly classified samples. To identify such classes, we study in this work the Class False Positive Score. We further argue that this perspective provides interesting insights into the model behavior and potentially allows to improve our understanding of the model’s latent space and vulnerability.
@@ 2.1 Evaluatiuon Metrics
A higher CFPS for a class indicates that it is more susceptible to being mistakenly assigned to samples from other classes by the model. The classes that are most likely mistaken as other classes have a high chance of manipulation by attackers, which impacts the overall reliability and security of our classification model. This enables us to focus on improving the robustness of these vulnerable classes.
@@ 3 Experiments
Figure 1 illustrates the class-wise accuracy of the aforementioned architectures when evaluated on clean validation samples from the CIFAR10 dataset. The overall accuracy of each model allows us to categorize classes into two groups: strong classes and weak classes. This categorization is based on whether a class exhibits a class-wise accuracy above or below the average overall accuracy of the model. Notably, we observe that classes C3, C5, and C6, corresponding to ’bird’, ’deer’, and ’dog’ respectively, fall into the category of weak classes. This determination is made due to their relatively lower accuracy compared to the other classes across different robust models. It is important to emphasize that this pattern remains consistent regardless of the specific architecture employed for training the robust models. Nevertheless, it is also not possible to conclude from this type of class-wise accuracy evaluation that these weak classes are more susceptible to attacks than strong classes when targeting the attack on specific classes.
In the context of class-wise robust analysis, previous research has commonly identified the weak robust classes [26, 3]. These determinations were often made by calculating the class-wise accuracy deviations with respect to the overall model accuracy or strong class accuracy but they failed to see a common pattern of weak classes under the influence of common corruptions and attacks. However, such approaches may introduce potential biases, as the weakness of a class could be influenced by the overall model accuracy or the highest strong class accuracy. To ensure fairness and impartiality in our evaluation, we evaluate a metric called the Class False Positive Score, shortly CFPS. This metric focuses on model misclassifications among the classes independently of overall accuracy, enabling a comprehensive analysis of class-to-class biases exhibited by the models.
@@
We assess the consistency of class-wise robust classification accuracies and CFPSs with commonly corrupted sample types on the CIFAR10C dataset [14]. Figure 3 presents class-wise robust accuracies and CFPSs of the adversarially trained aforementioned models across various corruption types. Interestingly, weak classes still consistently exhibit the lowest robust classification accuracy even with the inclusion of common corruptions, while C4 (’cat’) still maintains the highest CFPS. i.e., the class vulnerabilities to misclassifications remained constant after the addition of common corruptions but the magnitude of vulnerability varies.
Which classes are more vulnerable to adversarial attacks, weak or highly misclassified? Experiments have revealed that the two indicators of the class-wise performance of the model (accuracy and CFPS) point to the distinct properties of the classes. A crucial question here is whether weak classes based on the least class-wise accuracy or those that are mostly misclassified as others are more susceptible to adversarial attacks. Therefore, we further investigate the influence of adversarial attacks [21, 7] on class-wise robustness and also evaluate the most likely targetable class under the influence of attacks. We consider PGD attack [21] using ResNet-50 [4] for this experimentation.
@@ Appendix B Similarities Between Different Model Predictions
The cosine similarity values are in the range of 0.76 to 0.83 across the aforementioned architectures, which depicts the similarity between the predictions of the aforementioned models is high.

### v2

@@ 1 Introduction
Convolutional neural networks (CNNs) have achieved widespread success in various vision applications including image classification [14, 19, 36], image segmentation[23], and object detection [30]. Nonetheless, the existence of adversarial examples [7, 22, 6, 1] and common corruptions [16] like blurring, zooming, or Gaussian noise, poses challenges in their real-world deployment. Extensive efforts have been devoted to defending against adversarial attacks and enhancing model generalization [35]. Adversarial training has emerged as a prominent defense technique to improve the robustness of classification models [2, 7]. Prior works have analyzed adversarial training from different perspectives including robust optimization [32], robust generalization [29, 10, 11], training strategy [37, 28, 33] and neural architecture [9, 8, 12, 24, 13, 20]. Howbeit, all these previous works have concentrated on improving the overall model robustness, neglecting the discrepancies in the robustness of individual classes. This imbalance in class-wise robustness can be exploited by attackers, who may target less robust classes. Therefore, a comprehensive understanding of adversarial training on class-wise robustness is crucial for improving the robustness of classification models in a meaningful way.
Recently a few studies have emphasized class-wise robustness disparity in adversarial training [31, 3]. However, their focus has been limited to comparing class-wise robust accuracy deviations to identify the vulnerable classes. While this is important, analyzing class-to-class biases is equally crucial for gaining insights into the latent space of robust models. Specifically, understanding which class labels are assigned erroneously or which classes are predominantly confused is essential. Therefore, we conduct a comprehensive study of class-wise robust accuracies with particular emphasis on false positives in class-wise misclassifications, to improve the understanding of class-wise biases.
@@ 2 Background: Network Evaluation Methods
When introducing new models for image classification tasks, network performance is typically evaluated in terms of accuracy [14, 36, 19]. Such evaluation is most important, as we do not need a network that classifies randomly and wrongly. Additionally, the evaluation of the network’s robustness received more popularity in the last few years. Hence, a variety of robustness measures has been proposed [18, 7]. On the one hand, common corruptions [16] have been introduced, which incorporate natural and system noise that can lead to misclassifications in the classification systems. On the other hand, adversarial attacks gained a lot of popularity to evaluate the network’s vulnerabilities. In consequence, a variety of attacks, e.g. [7, 22, 6] along with their defenses have been proposed [7, 37]. The robust accuracy of a model is thereby usually defined as the model’s accuracy under a specific adversarial attack or corruption. Thus, most of these studies focus on improving the overall robust accuracy of models under attacks or when facing corruptions. A few recent works further investigated the class biases in model accuracy and robustness, arguing for a fair training process that allows classifying all classes about equally well [34, 31, 26]. These works also showed that adversarial training seems to amplify class-wise biases in model accuracy. Yet, only little effort has been devoted to studying which classes pre-dominantly attract incorrectly classified samples. To identify such classes, we study in this work the Class False Positive Score. We further argue that this perspective provides interesting insights into the model behavior and potentially allows to improve our understanding of the model’s latent space and vulnerability.
@@ 2.1 Evaluatiuon Metrics
A higher CFPS for a class indicates that it is more susceptible to being mistakenly assigned to samples from other classes by the model. The classes that are most likely mistaken as other classes have a high chance of manipulation by attackers, which impacts the overall reliability and security of a classification model. This enables us to focus on improving the robustness of these vulnerable classes.
@@ 3 Experiments
Figure 1 illustrates the class-wise accuracy of the aforementioned architectures when evaluated on clean validation samples from the CIFAR10 dataset. The overall accuracy of each model allows us to categorize classes into two groups: strong classes and weak classes. This categorization is based on whether a class exhibits a class-wise accuracy above or below the average overall accuracy of the model. Notably, we observe that classes C3, C5, and C6, corresponding to ’bird’, ’deer’, and ’dog’ respectively, fall into the category of weak classes. This determination is made due to their relatively lower accuracy compared to the other classes across different robust models. It is important to emphasize that this pattern remains consistent regardless of the specific architecture employed for training the robust models. However, class-wise accuracy evaluation alone does not allow us to conclude that weaker classes are the primary source of confusion in model classification. Some classes may exhibit high accuracy yet still be frequently misclassified as other classes. Attackers can leverage this information to enhance the confusion in classifiers.
In the context of class-wise robust analysis, previous research has commonly identified the weak robust classes [31, 3]. These determinations were often made by calculating the class-wise accuracy deviations with respect to the overall model accuracy or strong class accuracy but they failed to see a common pattern of weak classes under the influence of common corruptions and attacks. Moreover, such approaches may introduce biases, as the perceived weakness of a class could be affected by overall model accuracy or the performance of the strongest class. To ensure fairness and impartiality in our evaluation, we evaluate a metric called the Class False Positive Score, shortly CFPS. This metric focuses on model misclassifications among the classes independently of overall accuracy, enabling a comprehensive analysis of class-to-class biases exhibited by the models.
@@
We assess the consistency of class-wise robust classification accuracies and CFPSs with commonly corrupted sample types on the CIFAR10-C dataset [17]. Figure 3 presents class-wise robust accuracies and CFPSs of the adversarially trained aforementioned models across various corruption types. Interestingly, weak classes still consistently exhibit the lowest robust classification accuracy even with the inclusion of common corruptions, while C4 (’cat’) still maintains the highest CFPS. i.e., the class vulnerabilities to misclassifications remained constant after the addition of common corruptions but the magnitude of vulnerability varies.
Which classes are more vulnerable to adversarial attacks, weak or highly misclassified? Experiments have revealed that the two indicators of the class-wise performance of the model (accuracy and CFPS) point to the distinct properties of the classes. A crucial question here is whether weak classes based on the least class-wise accuracy or those that are mostly misclassified as others are more susceptible to adversarial attacks. Therefore, we further investigate the influence of adversarial attacks [25, 7] on class-wise robustness and also evaluate the most likely targetable class under the influence of attacks. We consider PGD attack [25] using ResNet-50 [4] for this experiment.
@@ Appendix B Similarities Between Different Model Predictions
The cosine similarity values are in the range of 0.76 to 0.83 across the aforementioned architectures, which indicates that the similarity between the predictions of the aforementioned models is high.